OpenStudy (anonymous):

PLEASE HELP. WILL GIVE MEDAL AND BECOME FAN FOR EFFORT! Joshua is the place kicker for his college football team. Last season he kicked 42 times and never missed. Each field goal scored 3 points, and each extra point scored 1 point, for a total of 86 points. How many field goals did Joshua kick? A)19 B)20 C)22 D)23

OpenStudy (anonymous):

f = field goals p = point after f + p = 42 (Total number of kicks) 3f + p = 86 (Total number of points) Solve for p in the first equation p = 42 - f Now substitute that for p in the second equation 3f + (42 - f) = 86 Now just solve this equation for f 2f + 42 = 86 f = 22 Now we know f = 22, we can solve for p using either of the original equations 22 + p = 42 p = 20 Field goals = 22, Point after = 20

OpenStudy (anonymous):

Thank you lots!

OpenStudy (anonymous):

could i get a fan and medal:)

OpenStudy (anonymous):

I think I already gave you both. lol. :)